Last year the price of a color TV was Tk. 12,000 less than that of computer. This year, the price of TV was increased by 20% and computer was increased by Tk. 1,000 only. Now if you need Tk. 7,400 more to buy a TV and a computer, what is the current price of a computer?
Answer: 45000
I know the answer but i need the detailed solution.

Answer by stanbon(75887)

Last year the price of a color TV was Tk. 12,000 less than that of computer. This year, the price of TV was increased by 20% and computer was increased by Tk. 1,000 only. Now if you need Tk. 7,400 more to buy a TV and a computer, what is the current price of a computer?
-----
Let last year's price for a color TV be "x":
This years's price for a color TV is 1.2x:
----
Let last years price of a computer be x+12000
This year's price for a computer is x+13000
------
Equations:
Increased cost of TV and computer = 7400
0.2x + 1000 = 7,400
0.2x = 6,400
x = 64000/2
x = 32000 (last year's cost of color TV)
---
Last years cost of computer = x+12000 = 44,000
--------
Answer: This year's cost of computer = 44,000+1000 = 45000
==========
Cheers,
Stan H.
============